Calculate the sound intensity level  in decibels for a sound wave traveling in air and having a pressure amplitude of  =0.656 Pa.


1
Expert's answer
2021-06-02T13:11:21-0400

"Intensity I = \frac{P^2}{2 * d * v}"


Where P = 0.656 pa (Pressure in pascal)

d = Air Density = 1.29 kg/ cubic meter

v = Speed of sound in air = 331 m/sec.


"Therefore, Intensity I = \frac{0.656 * 0.656}{2 * 1.29 * 331} = 0.000504 W /sq. meter"


"Now, db = 10 * log \frac{I}{I0}"

I0 = 1

I = 0.000504

"Hence, db = 10 * log \frac{0.000504}{1} = 87 db"


Final Answer = 87 db